OpenStudy (anonymous):

Please help... Need answers! Will Fan! In the below system, solve for y in the first equation. 2x + y = 11 5x - 2y = -1 AND What is the value of y in the solution to the following system of equations? 5x - 3y = -3 2x - 6y = -6 Please Help!

OpenStudy (john_es):

Do you know the reduction or substitution methods?

OpenStudy (john_es):

For example, in the case of substitution you can solve the first equation for y \[y=11-2x\]And substitute this into the second equation, \[5x-2(11-2x)=-1\] And solve for x.

OpenStudy (john_es):

The same can be applied to the second system of equations you posted.

OpenStudy (anonymous):

I know of nether methods. I am so lost.

OpenStudy (john_es):

Can you solve the system in the form I left to you?

OpenStudy (anonymous):

5x−2(11−2x)=−1?

OpenStudy (anonymous):

Also, here are the answer choices: question 1: 2x + 11 -2x + 11 2x - 11 -2x - 11 question 2: -1 zero 1 2

OpenStudy (john_es):

Well, XD I think I gave you the answer in the first post. \[y=-2x+11\]

OpenStudy (john_es):

For question 2, you should solve the system by the substitution method.

OpenStudy (john_es):

Although it is easier if you use reduction.

OpenStudy (john_es):

I would rewrite the second system after divide by 2 the second equation, 5x - 3y = -3 x - 3y = -3 And then I would solve for x in the second, x=-3+3y And substitute in the first, 5(-3+3y)-3y=-3 And solve for y.

OpenStudy (john_es):

Do you understand it?

OpenStudy (john_es):

It seems that y=1, but better try to find the solution by yourself.

OpenStudy (anonymous):

Ok, I am just having trouble with the first equation now. I got 1 for the 2nd problem, correct?

OpenStudy (anonymous):

@John_es

OpenStudy (john_es):

Sorry, I was out. Yes, you was correct.